### v1

# Long relators in groups generated by two parabolic elements

Abstract.
We find a family of groups generated by a pair of parabolic elements in which every relator must admit a long subword of a specific form. In particular, this collection contains groups in which the number of syllables of any relator is arbitrarily large. This suggests that the existing methods for finding non-free groups with rational parabolic generators may be inadequate in this case, as they depend on the presence of relators with few syllables. Our results rely on two variants of the ping-pong lemma that we develop, applicable to groups that are possibly non-free. These variants aim to isolate the group elements responsible for the failure of the classical ping-pong lemma.
1. Introduction
1.1. Background and main results
Determining which pairs of parabolic elements of $\operatorname{SL}_{2}(\mathbb{C})$ (namely, elements with trace ±$\pm 2$) generate a free subgroup is a long-standing problem. For $\lambda,\mu\in\mathbb{C}$, denote throughout this paper
(1.1) $A_{\lambda}=\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix},\quad B_{\mu}=\begin{pmatrix}1&0\\ \mu&1\end{pmatrix}\qquad\text{and}\qquad G_{\lambda,\mu}=\langle A_{\lambda},B_{\mu}\rangle,\quad G_{\lambda}=\langle A_{\lambda},B_{\lambda}\rangle=G_{\lambda,\lambda}.$
Any two non-commuting parabolic matrices are simultaneously conjugate to ±$\pm A_{\lambda}$ and ±$\pm B_{\lambda}$ for some $\lambda\neq 0$ (see for example, [12, Appendix III]), and hence the problem reduces to the study of $G_{\lambda}$ (and we also note that $G_{\lambda}$ is free if and only if its image in $\operatorname{PSL}_{2}(\mathbb{C})$ is free). However, it will still be convenient later to replace the group $G_{\lambda}$ with a conjugate group, using the fact that when $\lambda_{1}\mu_{1}=\lambda_{2}\mu_{2}$, $G_{\lambda_{1},\mu_{1}}$ and $G_{\lambda_{2},\mu_{2}}$ are conjugate [12, Appendix III]. Let us call $\lambda\in\mathbb{C}$ free if $G_{\lambda}$ is free, and non-free otherwise. Every transcendental number is free [9], and Sanov famously proved that $2$ is free, while also characterizing the elements of $G_{2}$ [23]. Sanov’s result was later generalized by Brenner, who showed that every $\lambda\geq 2$ is free [6]. In [7, 18, 13], the well known ping-pong lemma was used to find large domains of free complex numbers.
Non-free numbers are abundant as well: they are dense in the unit ball in $\mathbb{C}$ [18] and real non-free numbers are dense in the interval $(-2,2)$ [7, 2]. Among non-free numbers, the rational ones are of particular interest: throughout the years, many non-free rationals contained in $\mathbb{Q}\cap(-2,2)$ have been found, whereas we are not aware of a single free number in this subset; it is conjectured that all $\lambda\in\mathbb{Q}\cap(-2,2)$ are non-free [16]. A key observation made by Lyndon and Ullman, in an effort to find such rationals, is the following: consider the word $w=A_{\lambda}^{n_{1}}B_{\lambda}^{n_{2}}\cdots A_{\lambda}^{n_{k}}$ where $n_{1},\dots,n_{k}\neq 0$. We call each $A_{\lambda}^{n_{2i+1}}$ and $B_{\lambda}^{n_{2i}}$ a syllable of $w$ (see Definition 1.1). Denote the upper right entry of the matrix $w$ by $p_{w}(\lambda)$. Thinking of $\lambda$ as a variable, $p_{w}(\lambda)$ turns out to be a polynomial, whose coefficients depend on the integers $n_{1},\dots,n_{k}$ and whose degree is precisely the number of syllables, $k$. Moreover, finding some word $w$ as above such that $p_{w}(\lambda)=0$ is equivalent to showing that $\lambda$ is non-free (in the sense defined earlier). This allows to find many non-free rationals by analyzing a corresponding infinite family of polynomials of low degree. Using only $k=3$, Lyndon and Ullman proved that for every positive integer $n$, the number $\frac{n}{n^{2}+1}$ is non-free [18], and Beardon managed to improve this by showing that $\frac{n}{jn^{2}\pm 1}$ is non-free for all positive integers $n$ and $j$ (with $nj\geq 2)$ [2]. Variants of this method were subsequently developed by numerous authors, including [25, 1, 16, 24] among others, to obtain more extensive families of non-free rationals. While this has become the most prominent method, several other approaches have also emerged, sometimes being employed in conjunction with the former. Linear recurrence sequences were used in [18, 5], and in the more recent paper [8] non-free rationals $\frac{p}{q}$ were found by showing that $G_{\frac{p}{q}}$ has finite index in $G_{\frac{1}{q}}$.
All of these techniques seem to share a common drawback: they are effective only when relators with few syllables are present in $G_{\frac{p}{q}}$; this should not be surprising, considering that as mentioned earlier, such relators correspond to low degree polynomials. For this reason, Beardon [2, p. 532] and the authors of [8, p. 53] have doubted that these techniques could be used to settle the conjecture for all $\mathbb{Q}\cap(-2,2)$, believing that relators in $G_{\lambda}$ tend to consist of more syllables as $\lambda\to 2$ (although the terminology in these papers is somewhat different). The motivation for the current paper has been to affirm their prediction, thereby showing the limitations of the currently available methods of finding non-free numbers in $\mathbb{Q}\cap(-2,2)$. Before presenting our results, let us introduce the following definitions and notation.
Definition 1.1.
Let $S$ be a finite set and $w$ a word composed of letters from $S$. Let $w=s_{1}^{n_{1}}\cdots s_{k}^{n_{k}}$ be its reduced form, i.e., $s_{i}\in S$, $s_{i}\neq s_{i+1}^{\pm 1}$ and $n_{i}\in\mathbb{Z}\setminus\{0\}$ for all $i$. The syllables of $w$ are the subwords $s_{1}^{n_{1}},\dots,s_{k}^{n_{k}}$. We denote by $\operatorname{syl}_{S}(w)$ the number of syllables in $w$ (namely, $k$). Now let $G$ be a group generated by $S\subset G\setminus\{1\}$, then we define
$\sigma_{S}(G)=\min\{\operatorname{syl}_{S}(w):w\text{ is a nonempty word composed of letters from }S\text{ and }w=1\text{ in }G\}.$
In other words, $\sigma_{S}(G)$ is the minimal number of syllables in a relator; throughout this paper, every mention of a relator refers to a nontrivial relator, namely, a nonempty word in the free group. If $S$ freely generates $G$, we define $\sigma_{S}(G)=\infty$. When $S$ is obvious from the context, we omit it and simply write $\operatorname{syl}(w)$ and $\sigma(G)$.
The following is our first main result. As mentioned, it shows that the number of syllables in relators of $G_{\lambda,2}$ increases as $\lambda\to 2$, but moreover, that every relator must admit a long subword of a specific form. It also includes a similar result for the groups $G_{\lambda,i}$ with $\lambda\to 2$.
Theorem 1.2.
Let $\mu\in\{2,i\}$. Then for every $N\in\mathbb{N}$ there exists a neighborhood $U\subset\mathbb{C}$ of $2$ such that for every $\lambda\in U$, every relator of $G_{\lambda,\mu}$ has a subword $w$ of the form:
  1. (1)
    if $\mu=2$: $w=(A_{\lambda}B_{2}^{-1})^{k}$ or $w=(A_{\lambda}^{-1}B_{2})^{k}$ where $k\geq N$,
  2. (2)
    if $\mu=i$: $w=A_{\lambda}^{n_{1}}B_{i}^{n_{2}}\cdots A_{\lambda}^{n_{k-1}}B_{i}^{n_{k}}$ where $k\geq N$ and $n_{j}\in\{\pm 1\}$.
In particular, $\sigma(G_{\lambda,\mu})\to\infty$ as $\lambda\to 2$. Furthermore, a lower bound on the length of $w$ can be computed effectively. More precisely, let $\lambda_{n}\to 2$, then there are effectively computable constants $k_{n}\in\mathbb{N}$ such that $k_{n}\to\infty$, and any relator of $G_{\lambda_{n},\mu}$ has a subword $w_{n}$ of the above form with $\operatorname{syl}(w_{n})\geq k_{n}$.
This result does not depend on the specific chosen normalization of the groups $G_{\lambda,\mu}$, as the following corollary indicates.
Corollary 1.3.
Let $(\lambda_{n})_{n=1}^{\infty},(\mu_{n})_{n=1}^{\infty}\subset\mathbb{C}$ and suppose that $\lambda_{n}\cdot\mu_{n}\to\ell$ where {±$\ell\in\{\pm 4,\pm 2i\}$. Then for every $K\in\mathbb{N}$ there exists $N\in\mathbb{N}$ such that for every $n\geq N$, every relator of $G_{\lambda_{n},\mu_{n}}$ has a subword $w$ of the form:
  1. (1)
    if =±$\ell=\pm 4$: $w=(A_{\lambda_{n}}\,B_{\mu_{n}}^{-1})^{k}$ or $w=(A_{\lambda_{n}}^{-1}\,B_{\mu_{n}})^{k}$ where $k\geq K$,
  2. (2)
    if =±$\ell=\pm 2i$: $w=A_{\lambda_{n}}^{n_{1}}\,B_{\mu_{n}}^{n_{2}}\cdots A_{\lambda_{n}}^{n_{k-1}}\,B_{\mu_{n}}^{n_{k}}$ where $k\geq K$ and $n_{j}\in\{\pm 1\}$.
In particular, $\sigma(G_{\lambda_{n},\mu_{n}})\underset{n\to\infty}{\longrightarrow}\infty$.
In the case of real parameters, we are able to give an explicit lower bound on the length of this subword.
Theorem 1.4.
There exists a constant $C>0$ such that for every real $21.8\leq\lambda<$ 2, every (nontrivial) relator of $G_{\lambda,2}$ admits a subword of the form $(A_{\lambda}B_{2}^{-1})^{k}$ or $(A_{\lambda}^{-1}B_{2})^{k}$, where $k\geq\frac{C}{\sqrt{2-\lambda}}$. In particular, $\sigma(G_{\lambda,2})\geq\frac{2C}{\sqrt{2-\lambda}}$.
This asymptotic lower bound is sharp; see Remark 3.3 for the details.
As $\lambda\to 2$, the generators $A_{\lambda},B_{2}$ of $G_{\lambda,2}$ converge to $A_{2},B_{2}$ which generate the free group of rank $2$. Theorem 1.2 then shows that $\sigma(G_{\lambda,2})\to\sigma(G_{2,2})=\infty$. We stress that in general, $\sigma$ is not necessarily semicontinuous, as the following example demonstrates.
Example 1.5.
For $\lambda\in\mathbb{R}$ let $C_{\lambda}=\begin{pmatrix}1&2\\ \frac{\lambda-1}{2}&\lambda\end{pmatrix}\in\operatorname{SL}_{2}(\mathbb{R})$. Then $C_{1}=A_{2}$, and since $A_{2},B_{2}$ freely generate $G_{2}$, $\sigma(\langle C_{1},B_{2}\rangle)=\sigma(G_{2})=\infty$. However, let us show that $\liminf_{\lambda\to 1}\sigma(\langle C_{\lambda},B_{2}\rangle)=1$. We recall that every element in $\operatorname{SL}_{2}(\mathbb{R})$ can be classified as either elliptic, parabolic or hyperbolic according to its trace. An element $g\in\operatorname{SL}_{2}(\mathbb{R})$ with $\lvert\operatorname{tr}g\rvert=\lvert 2\cos\theta\rvert<2$ is elliptic, and acts on the hyperbolic half-plane $\mathbb{H}^{2}$ as a (hyperbolic) rotation of angle ±$\pm 2\theta$; for more details, see for example [3, 14, 4]. Thus, if we define $\lambda_{n}=2\cos(\pi/n)-1$ for every $n\in\mathbb{N}$, then we obtain a sequence $\lambda_{n}\to 1$ such that $C_{\lambda_{n}}$ is a rotation of angle $2\pi/n$, hence $C_{\lambda_{n}}^{2n}=1$ and the assertion follows.
These results are based on two variations of the ping-pong lemma that we introduce (Props. 2.5 and 2.8), which can be applied to a group regardless of whether it is free or not. We will apply them to the actions of the group $G_{\lambda,\mu}$ on the real and complex projective lines. The difficulty when dealing with syllables lies in the fact that a single syllable may consist of an arbitrarily large power, thereby making it harder to control the behaviour of (the non-equicontinuous family) $\{A_{\lambda}^{n}:n\neq 0\}$ as $\lambda$ varies. However, one key observation is that the majority of group elements act nicely on those spaces. The failure of the ping-pong lemma can only be attributed to very specific elements, which must be present in every relator. This observation will permit us to confine our attention to only finitely many group elements and use continuity arguments. In Section 2 we will develop these variations of the ping-pong lemma and define relevant notions. This will be done in general settings, and in Section 3 we will revisit the groups $G_{\lambda,\mu}$ and prove the main results.
1.2. Discussion and questions
We conclude this section with a few remarks, questions that came up during this work and possible future research directions.
  1. (1)
[…]
  2. (2)
    In addition to the freeness of $G_{1,\mu}$ (for $\mu\in\{\pm 4,\pm 2i\}$), its discreteness allows us to use our ping-pong argument. Therefore, the following natural question arises: let
    =$\mathcal{R}^{\prime}=\{0\neq\mu\in\mathbb{C}:G_{1,\mu}\text{ is free and discrete}\}.$
    For what values $\mu\in\partial\mathcal{R}^{\prime}$ does Theorem 1.2 hold? It turns out that $\partial\mathcal{R}^{\prime}$ is a Jordan curve, and furthermore, $\mathcal{R}^{\prime}$ is precisely the closure of the well studied Riley slice [21] (for more information about the Riley slice, see for example [15, 17]). It seems to be reasonable that our results could be extended to other values $\mu\in\partial\mathcal{R}^{\prime}$, however, some new complications arise. The primary utility of the specific values of $\mu$ is in selecting $X_{1}$ and $X_{2}$, the corresponding subsets for the ping-pong argument. However, such subsets may not exist for all $\mu\in\partial\mathcal{R}^{\prime}$, as the ordinary set $\Omega(G_{1,\mu})$ might be empty [19]. Nevertheless, the ordinary set is nonempty for a dense collection of $\mu\in\partial\mathcal{R}^{\prime}$ [20]; groups $G_{1,\mu}$ associated with such $\mu$ are called cusp groups. Let us also note that all of the cusp groups with $\mu\in\partial\mathcal{R}^{\prime}\setminus\{\pm 4,\pm 2i\}$ are non-classical, meaning that the Jordan curves defining the boundary of the corresponding subsets $X_{1}$ and $X_{2}$ will neither be circles nor lines (see [11, Theorem 6.18] and also [10, Figure 3]).
  3. (3)
    We return to the normalization $G_{\lambda,2}$ which will be used throughout most of the paper. For every $n\in\mathbb{N}$ define $V_{n}=\{\lambda\in\mathbb{C}:\sigma(G_{\lambda,2})\geq n\}$ and $U_{n}=V_{n}\cap\mathbb{R}$. Theorem 1.2 asserts that for every $n$, $V_{n}$ contains complex open neighborhoods of ±$\pm 2$ and ±$\pm i$ and $U_{n}$ contains real open neighborhoods of ±$\pm 2$. It would be interesting to study the properties of these sets further; in particular, whether they are open or contain a dense subset of rationals. As previously discussed, many papers have studied groups that admit short relators (syllables-wise), and such groups correspond to values $\lambda\in\mathbb{C}\setminus V_{n}$ and $\lambda\in\mathbb{R}\setminus U_{n}$ for small $n$. Moreover, these values of $\lambda$ admit infinitely many accumulation points [18, 2, 25, 24], but this does not rule out the possibility that $\mathbb{C}\setminus V_{n}$ and $\mathbb{R}\setminus U_{n}$ are, for example, nowhere dense.
  4. (4)
    It would be interesting to know if our results extend to a larger collection of matrices which converge to $A_{2}$. For instance, can we replace $A_{\lambda}$ with some hyperbolic element such as
    $A_{\lambda}^{\prime}=\begin{pmatrix}1+2\lambda&2\\ \lambda&1\end{pmatrix}$
    (with $\lambda>0$)? Of course, this would only be interesting if not all of the groups $\langle A_{\lambda}^{\prime},B_{2}\rangle$ are free.
  5. (5)
    Let $w$ be a relator of $G_{\lambda,\mu}$, where $\mu\in\{2,i\}$. Theorem 1.2 tells us that $w$ has a long subword $w^{\prime}$ of a particular form, however, it gives us no information on whether $w^{\prime}$ constitutes a large chunk of $w$, namely, on $\operatorname{syl}(w^{\prime})/\operatorname{syl}(w)$ or $\operatorname{len}(w^{\prime})/\operatorname{len}(w)$ (recall that the length of a reduced word $w$, denoted $\operatorname{len}(w)$, is its number of letters).
2. The ping-pong lemma
In this section, we aim to formulate and prove two variations of the ping-pong lemma: one for a single group and one for a family of groups. We begin by introducing a particular decomposition of words in the free group, which will be employed in the proofs. Let $F=F_{S}$ be the free group generated by some set $S$ with $2\leq\lvert S\rvert<\infty$.
Definition 2.1.
Let $w=s_{1}^{n_{1}}\cdots s_{k}^{n_{k}}\in F$ be a reduced word with $s_{i}\in S$. An exact subword of $w$ is any nonempty subword composed of syllables, i.e., of the form $s_{i}^{n_{i}}\cdots s_{j}^{n_{j}}$ where $i\leq j$.
Definition 2.2.
Let $W\subset F$ be any subset. Thinking of $W$ as forbidden words, we define $\mathcal{F}(W)\subset F$ to be the collection of all nonempty words avoiding $W$, namely
$\mathcal{F}(W)=\{w\in F:w\text{ is nonempty and no exact subword of }w\text{ belongs to }W\}.$
Note that $W\cap\mathcal{F}(W)=\emptyset$. An $\mathcal{F}(W)$-decomposition of a word $w\in F$ is a decomposition $w=w_{1}\cdots w_{n}$ into exact subwords, such that for every $i$:
  1. (1)
    $w_{i}\in W\cup\mathcal{F}(W)$,
  2. (2)
    no two consecutive subwords belong to $\mathcal{F}(W)$, i.e., if $w_{i}\in\mathcal{F}(W)$ then $w_{i+1}\in W$.
Lemma 2.3.
Let $W\subset F$ be any subset. Then every nonempty word $w\in F$ has an $\mathcal{F}(W)$-decomposition.
Proof.
Let $w\in F$ be nonempty. Consider the collection of all decompositions of $w$ into exact subwords $w=w_{1}\cdots w_{n}$ such that |$\big{\lvert}\{1\leq i\leq n:w_{i}\in W\}\big{\rvert}=k$, where $k\in\mathbb{N}$ is the maximal possible [^1] Note that the empty word is not considered an exact subword, and hence $w$ has only finitely many decompositions.. Among this collection, choose a decomposition $w=w_{1}\cdots w_{n}$ such that |$\big{\lvert}\{1\leq i\leq n:w_{i}\in\mathcal{F}(W)\}\big{\rvert}$ is the minimal possible; we claim that this is an $\mathcal{F}(W)$-decomposition. Indeed, if for some $i$, $w_{i}\notin\mathcal{F}(W)$, then there exists a decomposition $w_{i}=v_{1}\,u\,v_{2}$, where $u\in W$ is an exact subword of $w$, and each $v_{j}$ is either an exact subword or the empty word. By the maximality of $k$, the decomposition $w_{1}\cdots w_{i-1}\,v_{1}\,u\,v_{2}\,w_{i+1}\cdots w_{n}$ contains no more subwords from $W$ than the former decomposition $w_{1}\cdots w_{n}$. This implies that $w_{i}\in W$, yielding property $(1)$. Next, suppose for contradiction that $w_{i},w_{i+1}\in\mathcal{F}(W)$. Since this decomposition has the minimal possible number of subwords in $\mathcal{F}(W)$ (among the specified collection), we conclude that $w_{i}\,w_{i+1}\notin\mathcal{F}(W)$. This means that we can decompose $w_{i}\,w_{i+1}=v_{1}\,u\,v_{2}$ where once again $u\in W$ is an exact subword and $v_{1},v_{2}$ are exact subwords or empty. Replacing $w_{i}\,w_{i+1}$ with $v_{1}\,u\,v_{2}$, we obtain another decomposition of $w$ which contains more than $k$ subwords from $W$, contradicting the maximality of $k$. ∎
Let $H=\langle a,b\rangle$ be a group acting on a set $X$. Let $F_{2}$ be the free group generated by the symbols $\overline{a}$ and $\overline{b}$. From now on it will be important to distinguish between the free group elements $\overline{w}=\overline{w}(\overline{a},\overline{b})\in F_{2}$ and the corresponding element $w=\overline{w}(a,b)\in H$.
For $x,y\in\{a,b\}$ we denote by $F_{2}^{x,y}$ the collection of all words in $F_{2}$ whose first letter is $\overline{x}$ and last letter is $\overline{y}$.
Definition 2.4.
For any four subsets $C_{1},C_{2},D_{1},D_{2}\subset X$ define $W(C_{1},C_{2},D_{1},D_{2})\subset F_{2}$ to be the collection of all $\overline{w}\in F_{2}$ such that
(2.1) $\begin{cases}w(C_{1})\subset D_{2}&\text{if }\overline{w}\in F_{2}^{a,a}\\ w(C_{2})\subset D_{2}&\text{if }\overline{w}\in F_{2}^{a,b}\\ w(C_{1})\subset D_{1}&\text{if }\overline{w}\in F_{2}^{b,a}\\ w(C_{2})\subset D_{1}&\text{if }\overline{w}\in F_{2}^{b,b}.\end{cases}$
For $i=1,2$, let $K_{i}\subset X_{i}\subset X$ be some subsets. We will be interested in the two sets $W(K_{1},K_{2},X_{1},X_{2})$ and $W(X_{1},X_{2},K_{1},K_{2})$, and it will be helpful to think about them in the following context: for any $n\neq 0$, $a^{n}\in H$ ($b^{n}\in H$) is a ping-pong player whose goal is to pass elements of $X_{1}$ into $X_{2}$ (from $X_{2}$ into $X_{1}$). Every word can be thought of as an iteration of ping-pong shots, according to its syllables. For example, the goal of $b^{-1}a^{2}$ is to pass $X_{1}$ back into $X_{1}$. Since we will deal with non-free groups, some words will fail to achieve this. The former set is the collection of words that still manage to pass “easier” elements, namely, elements of $K_{1}$ and $K_{2}$, even if they possibly fail to pass some other elements of $X_{1}$ and $X_{2}$. The latter set consists of the outstanding players who not only manage to make the passes from $X_{1}$ and $X_{2}$, but to put them inside $K_{1}$ and $K_{2}$.
The following is our first variant of the ping-pong lemma. Essentially, it states that every relator of the group, whose first and last letters are identical, contains a subword $w$ that fails to pass elements from $K_{1}$ and $K_{2}$ into $X_{1}$ and $X_{2}$, and furthermore, no subword of $w$ succeeds in passing $X_{1}$ and $X_{2}$ into $K_{1}$ and $K_{2}$. This will prove useful in the next section, as we will see that with a suitable choice of $X_{1},X_{2},K_{1},K_{2}$ there are relatively few words $w$ exhibiting these properties.
Proposition 2.5.
Let $H=\langle a,b\rangle$ be a group acting on a set $X$. Let $F_{2}$ be the free group generated by the symbols $\overline{a}$ and $\overline{b}$. Let $X_{1},X_{2}\subset X$ be disjoint, and $K_{i}\subset X_{i}$ be nonempty for $i=1,2$. Denote
$W_{X\to K}=W(X_{1},X_{2},K_{1},K_{2})\subset F_{2}\qquad\text{and}\qquad W_{K\to X}=W(K_{1},K_{2},X_{1},X_{2})\subset F_{2}.$
Then every relator of $H$, whose first and last letters are identical, has an exact subword that belongs to $\mathcal{F}(W_{X\to K})\setminus W_{K\to X}$.
If in addition, there exists $N\in\mathbb{N}$ such that $\overline{a}^{k}\in W_{X\to K}$ for every $\lvert k\rvert\geq N$, then the assumption on the first and last letters of the relator can be dropped.
We will need the following lemma.
Lemma 2.6.
Let $\overline{w}\in F_{2}$ and $\overline{w}=\overline{w}_{1}\cdots\overline{w}_{n}$ be a decomposition into exact subwords that alternate between $W_{K\to X}$ and $W_{X\to K}$, i.e., either $\overline{w}_{2i}\in W_{K\to X}$ and $\overline{w}_{2i+1}\in W_{X\to K}$ or the other way around. Then $\overline{w}\in W_{K\to X}$.
Proof of Lemma 2.6.
Let us assume that $\overline{w}_{n}\in W_{K\to X}$, and that the first and last letters of $\overline{w}_{n}$, and also the first letter of $\overline{w}_{n-1}$, are $\overline{a}$; the proof of all other cases is similar. Since $\overline{w}_{n-1}$ and $\overline{w}_{n}$ are exact subwords, the last letter of $\overline{w}_{n-1}$ must differ from the first letter of $\overline{w}_{n}$, and therefore it must be $\overline{b}$. We obtain
$w_{n-1}w_{n}(K_{1})\subset w_{n-1}(X_{2})\subset K_{2}.$
Proceeding in the same manner and recalling that $K_{i}\subset X_{i}$ in case $n$ is even, we see that $\overline{w}\in W_{K\to X}$. ∎
Proof of Prop. 2.5.
Let $\overline{w}\in F_{2}$ be a nonempty word that has no exact subword from $\mathcal{F}(W_{X\to K})\setminus W_{K\to X}$, and let $\overline{w}=\overline{w}_{1}\cdots\overline{w}_{n}$ be some $\mathcal{F}(W_{X\to K})$-decomposition (Lemma 2.3). We will show by induction on |$\big{\lvert}\{i:\overline{w}_{i}\notin W_{X\to K}\}\big{\rvert}$ that $\overline{w}\in W_{K\to X}$. If $\overline{w}_{i}\in W_{X\to K}$ for all $i$, then $\overline{w}\in W_{K\to X}$; this can be seen as a simple case of Lemma 2.6 since $W_{X\to K}\subset W_{K\to X}$. Next, suppose that for some $i$, $\overline{w}_{i}\notin W_{X\to K}$. By the first property of the $\mathcal{F}(W_{X\to K})$-decomposition (Definition 2.2) $\overline{w}_{i}\in\mathcal{F}(W_{X\to K})$ and by our assumption on $\overline{w}$, $\overline{w}_{i}\in W_{K\to X}$. If $i=1$ let us define $\overline{w}_{i-1}$ to be the empty word, and otherwise, the second property of the $\mathcal{F}(W_{X\to K})$-decomposition implies that $\overline{w}_{i-1}\in W_{X\to K}$. Similarly, $\overline{w}_{i+1}$ is either empty if $i=n$ or belongs to $W_{X\to K}$ as well. If $i>2$, then $\overline{u}_{1}\coloneqq\overline{w}_{1}\cdots\overline{w}_{i-2}\in W_{K\to X}$ by the induction hypothesis and if $i\leq 2$ we let $\overline{u}_{1}$ be the empty word. Similarly, $\overline{u}_{2}\coloneqq\overline{w}_{i+2}\cdots\overline{w}_{n}\in W_{K\to X}$ if $i<n-1$ and $\overline{u}_{2}$ is empty otherwise. Decomposing $\overline{w}$ as $\overline{w}=\overline{u}_{1}\,\overline{w}_{i-1}\,\overline{w}_{i}\,\overline{w}_{i+1}\,\overline{u}_{2}$, we deduce from Lemma 2.6 that $\overline{w}\in W_{K\to X}$, completing the induction. In particular, if $\overline{w}\in F_{2}^{a,a}$ then $w(K_{1})\subset X_{2}$ which implies that $w\neq 1$, and a similar argument applies if $\overline{w}\in F_{2}^{b,b}$.
For the second part, let $\overline{w}$ be a relator of $H$. Choose $m$ such that the first and last syllables of $\overline{w}^{\prime}\coloneqq\overline{a}^{m}\,\overline{w}\,\overline{a}^{-m}$ are $\overline{a}^{k},\overline{a}^{k^{\prime}}$ for some $\lvert k\rvert,\lvert k^{\prime}\rvert\geq N$. Then $\overline{w}^{\prime}$ has an exact subword from $\mathcal{F}(W_{X\to K})\setminus W_{K\to X}$, and it must also be an exact subword of $\overline{w}$. ∎
Corollary 2.7.
Keeping the notation of Proposition 2.5, suppose that
$\{\overline{a}^{n},\overline{b}^{n}:\lvert n\rvert\geq 2\}\subset W_{X\to K},$
and denote
(2.2) $\mathcal{A}=\{\overline{w}\in F_{2}:\text{all syllables of }\overline{w}\text{are either }\overline{a},\overline{a}^{-1},\overline{b}\text{ or }\overline{b}^{-1}\}$
(e.g. $w=aba^{-1}b$). Then every relator of $H$ has an exact subword that belongs to $\mathcal{A}\setminus W_{K\to X}$.
If in addition
(2.3) $\{(\overline{a}\overline{b})^{\pm 1},(\overline{b}\overline{a})^{\pm 1}\}\subset W_{X\to K},$
Then every relator of $H$ has an exact subword that belongs to $\mathcal{B}\cup\mathcal{B}^{-1}\setminus W_{K\to X}$, where
(2.4) =$\mathcal{B}=\{(\overline{a}\overline{b}^{-1})^{k},\quad(\overline{b}^{-1}\overline{a})^{k},\quad(\overline{a}\overline{b}^{-1})^{k-1}\overline{a},\quad(\overline{b}^{-1}\overline{a})^{k-1}\overline{b}^{-1}:k>0\}$
and $\mathcal{B}^{-1}=\{\overline{w}^{-1}:\overline{w}\in\mathcal{B}\}$.
Proof.
In every word that admits no exact subword from $W_{X\to K}$, every letter $\overline{a}^{\pm 1}$ can be followed only by $\overline{b}^{\pm 1}$ and vice versa. Thus, $\mathcal{F}(W_{X\to K})\subset\mathcal{A}$. Under the additional assumption (2.3), in every word in $\mathcal{F}(W_{X\to K})$ each letter $\overline{a},\overline{a}^{-1},\overline{b},\overline{b}^{-1}$ can be followed only by the letter $\overline{b}^{-1},\overline{b},\overline{a}^{-1},\overline{a}$ respectively. Therefore in this case, $\mathcal{F}(W_{X\to K})\subset\mathcal{B}\cup\mathcal{B}^{-1}$. The corollary now follows from Proposition 2.5. ∎
Until now, we have focused solely on a single group generated by the elements $a,b$. We will now consider the scenario where $a,b$ vary. To be more precise, let $H$ be a topological group and $\Lambda$ a metric space. Then we let $a,b:\Lambda\to H$ be continuous functions. Given $\lambda\in\Lambda$, it will be convenient to denote $a_{\lambda}\coloneqq a(\lambda)$ and $b_{\lambda}\coloneqq b(\lambda)$. Suppose that $H$ acts continuously on a metric space $X$. For subsets $X_{1},X_{2},K_{1},K_{2}\subset X$, let us denote
$W_{X\to K}^{\lambda}=W(X_{1},X_{2},K_{1},K_{2})\qquad\text{and}\qquad W_{K\to X}^{\lambda}=W(K_{1},K_{2},X_{1},X_{2}).$
where the right-hand side of both is defined with respect to the group $H_{\lambda}\coloneqq\langle a_{\lambda},b_{\lambda}\rangle$ (see Definition 2.4). We recall that these are subsets of the free group $F_{2}$ generated by the symbols $\overline{a}$ and $\overline{b}$. Here is our second variant of the ping-pong lemma.
Proposition 2.8.
In the above settings, suppose that $X_{1}$ and $X_{2}$ are disjoint and open, and that $K_{i}\subset X_{i}$ are nonempty compact subsets for $i=1,2$. Suppose further that for some $\lambda_{0}\in\Lambda$ and every $n\neq 0$, $a_{\lambda_{0}}^{n}X_{1}\subset X_{2}$ and $b_{\lambda_{0}}^{n}X_{2}\subset X_{1}$. Then for every $N\in\mathbb{N}$ there exists a neighborhood $U\subset\Lambda$ of $\lambda_{0}$ with the following property: let $\lambda\in U$, and $\overline{w}$ a relator of $H_{\lambda}$ whose first and last letters are identical. Then $\overline{w}$ has an exact subword from $\mathcal{F}(W_{X\to K}^{\lambda})$ whose length is at least $N$. If in addition, there exists $M\in\mathbb{N}$ such that $\overline{a}^{k}\in W_{X\to K}^{\lambda}$ for every $\lvert k\rvert\geq M$ and $\lambda\in\Lambda$, then the assumption on the first and last letters of $\overline{w}$ can be dropped.
Proof.
First we recall that $\operatorname{len}(\overline{w})$ denotes the number of letters of the word $\overline{w}$, and that $F_{2}^{x,y}\subset F_{2}$ consists of the words that begin with the letter $\overline{x}$ and end with $\overline{y}$. Let $N\in\mathbb{N}$, and define $L=\{w\in F_{2}:\operatorname{len}(w)<N\}$. For $x,y\in\{a,b\}$, we denote $L^{x,y}=F_{2}^{x,y}\cap L$, and for $\lambda\in\Lambda$, $L_{\lambda}^{x,y}$ is the subset of corresponding elements in $H_{\lambda}$ (i.e., the image of $L^{x,y}$ under the homomorphism $\overline{a}\mapsto a_{\lambda},\overline{b}\mapsto b_{\lambda}$). Denote by $d$ the metric on $X$. Since $L$ is finite and $K_{1}$ is compact, the function
$h:\Lambda\to\mathbb{R},\qquad h(\lambda)=d(L^{a,a}_{\lambda}K_{1},X\setminus X_{2})$
(where $L^{a,a}_{\lambda}K_{1}\coloneqq\{gx:g\in L^{a,a}_{\lambda},x\in K_{1}\}$) is continuous. By our assumption on $\lambda_{0}$, $L^{a,a}_{\lambda_{0}}K_{1}\subset X_{2}$ and hence $h>0$ on some neighborhood $U\subset\Lambda$ of $\lambda_{0}$. In other words, $L^{a,a}_{\lambda}K_{1}\subset X_{2}$ for every $\lambda\in U$. Arguing analogously and taking smaller $U$ if necessary, we can see that for every $\lambda\in U$,
(2.5) $L^{a,a}_{\lambda}K_{1}\subset X_{2},\qquad L^{a,b}_{\lambda}K_{2}\subset X_{2},\qquad L^{b,a}_{\lambda}K_{1}\subset X_{1}\qquad\text{and}\qquad L^{b,b}_{\lambda}K_{2}\subset X_{1}.$
This means that for every $\lambda\in U$, $L\subset W_{K\to X}^{\lambda}$. In other words, all words in $\mathcal{F}(W_{X\to K}^{\lambda})\setminus W_{K\to X}^{\lambda}$ are of length at least $N$. We conclude by applying Proposition 2.5. ∎
Corollary 2.9.
Under the assumptions of the Proposition 2.8, suppose that there exists $M\in\mathbb{N}$ such that $\overline{a}^{k},\overline{b}^{k}\in W_{X\to K}^{\lambda}$ for every $\lvert k\rvert\geq M$ and $\lambda\in\Lambda$. Then $\sigma(H_{\lambda})\to\infty$ as $\lambda\to\lambda_{0}$ (cf. Definition 1.1).
Proof.
Let $\lambda_{n}\to\lambda_{0}$ and for every $n$ let $\overline{w}_{n}$ be a relator of $H_{\lambda_{n}}$. Then by Proposition 2.8, there is a sequence $(\overline{u}_{n})$ such that every $\overline{u}_{n}$ is an exact subword of $\overline{w}_{n}$ that belongs to $\mathcal{F}(W_{X\to K}^{\lambda_{n}})$ and $\operatorname{len}(\overline{u}_{n})\to\infty$. By our assumption, every syllable of $\overline{u}_{n}$ is of the form $a^{k}$ or $b^{k}$ where $\lvert k\rvert<M$. Since for any $C\in\mathbb{N}$, there are only finitely many such words with less than $C$ syllables, it follows that $\operatorname{syl}(\overline{w}_{n})\geq\operatorname{syl}(\overline{u}_{n})\underset{n\to\infty}{\longrightarrow}\infty$. Since $(\lambda_{n})$ and $(\overline{w}_{n})$ were arbitrary, $\sigma(H_{\lambda})\to\infty$. ∎
Corollary 2.10.
Under the assumptions of Proposition 2.8, suppose that for every $\lambda\in\Lambda$
(2.6) $\{\overline{a}^{n},\overline{b}^{n}:\lvert n\rvert\geq 2\}\subset W_{X\to K}^{\lambda}.$
Then for every $N\in\mathbb{N}$ there exists a neighborhood $U\subset\Lambda$ of $\lambda_{0}$ such that for every $\lambda\in U$, every relator of $H_{\lambda}$ has an exact subword of the form $\overline{a}^{n_{1}}\overline{b}^{n_{2}}\cdots\overline{a}^{n_{k-1}}\overline{b}^{n_{k}}$ where $k\geq N$ and $n_{j}\in\{\pm 1\}$.
If in addition,
(2.7) $\{(\overline{a}\overline{b})^{\pm 1},(\overline{b}\overline{a})^{\pm 1}\}\subset W_{X\to K}^{\lambda},$
then the form of the mentioned subword is $(\overline{a}\overline{b}^{-1})^{k^{\prime}}$ or $(\overline{a}^{-1}\overline{b})^{k^{\prime}}$ where $2k^{\prime}\geq N$.
Proof.
Let $U\subset\Lambda$ be the neighborhood of $\lambda_{0}$ given by Proposition 2.8. Let $\lambda\in U$ and let $w$ be a relator of $H_{\lambda}$. As was seen in the proof of Corollary 2.7, $\mathcal{F}(W_{X\to K}^{\lambda})\subset\mathcal{A}$ (where $\mathcal{A}$ is as in Equation (2.2)), and hence by Proposition 2.8, $w$ has an exact subword from $\mathcal{A}$ whose length is at least $N$. If this subword begins with the letter $b^{\pm 1}$, or ends with $a^{\pm 1}$, we can omit these letters to obtain a subword of the desired form. The second part follows similarly, using the fact that $\mathcal{F}(W_{X\to K}^{\lambda})\subset\mathcal{B}\cup\mathcal{B}^{-1}$ (where $\mathcal{B}$ is defined in Equation (2.4)). ∎
3. Proofs of the main results
In this section, we shift our focus back to the groups defined in the introduction, and prove all of the main results. Recall the notation $A_{\lambda},B_{\lambda}$ from Equation (1.1). We consider the standard action of $\operatorname{SL}_{2}(\mathbb{C})$ by Möbius transformations on the extended complex plane $^{}^\widehat{\mathbb{C}}\coloneqq\mathbb{C}\cup\{\infty\}$. In particular, $A_{\lambda}^{n}z=z+n\lambda$ and $B_{\lambda}^{n}z=\frac{z}{n\lambda z+1}$ for every $z\in\widehat{\mathbb{C}}$. We begin with a simple observation.
Lemma 3.1.
Let $\lambda,\mu\in\mathbb{C}$ and let $w=s_{1}^{n_{1}}\cdots s_{k}^{n_{k}}$ be a reduced word with $s_{i}\in\{A_{\lambda},B_{\mu}\}$. Then using the convention =$\infty$=$\infty$- ∞ = ∞, for every $z\in\widehat{\mathbb{C}}$,
$s_{1}^{-n_{1}}\cdots s_{k}^{-n_{k}}(z)=-w(-z).$
Proof.
The claim is readily checked for $w=A_{\lambda}^{n}$ or $w=B_{\mu}^{n}$, and hence follows by induction on $\operatorname{syl}(w)$. ∎
Proof of Theorem 1.2 for $\mu=2$.
Let $\Lambda=\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda>1.5\}$, and define $a,b:\Lambda\to\operatorname{SL}_{2}(\mathbb{C})$ by
(3.1) $a_{\lambda}\coloneqq a(\lambda)=\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix}=A_{\lambda},\qquad b=b(\lambda)=\begin{pmatrix}1&0\\ 2&1\end{pmatrix}=B_{2}$
(the function $b$ is constant). Define
(3.2) $X_{1}=\{z\in\mathbb{C}:\lvert z\rvert<1\}\subset\widehat{\mathbb{C}},\qquad X_{2}=\{z\in\mathbb{C}:\lvert z\rvert>1\}\cup\{\infty\}\subset\widehat{\mathbb{C}}.$
We note that for every $n\geq 1$ and $\lambda\in\Lambda$,
(3.3) $\operatorname{Re}(a_{\lambda}^{n}z)=\operatorname{Re}z+n\operatorname{Re}\lambda$
and also
It follows from these observations that for every $n\geq 1$, $a_{2}^{n}(X_{1})\subset X_{2}$ and $b^{n}(X_{2})\subset X_{1}$. By Lemma 3.1 and the symmetry (about 00) of $X_{1}$ and $X_{2}$, the same holds for $n<0$ as well, verifying the assumption of Proposition 2.8 for $\lambda_{0}=2$.
Next, we define
(3.5) $K_{1}=\{\lvert z\rvert\leq 1/2\},\qquad K_{2}=\{\lvert z\rvert\geq 5/4\}\cup\{\infty\},$
and turn to verify Equations (2.6) and (2.7). Let $\lambda\in\Lambda$. Once again by Lemma 3.1, it suffices to verify those inclusions for the words with positive powers. Let $n\geq 2$. By Equation (3.3), $a_{\lambda}^{n}X_{1}\subset K_{2}$ and by Equation (3.4), the leftmost point of $b^{n}X_{2}$ is positive and its rightmost point is bounded above by $1/3$, so $b^{n}X_{2}\subset K_{1}$. Equation (3.4) also implies that for $z\in X_{2}$, $\operatorname{Re}bz>1/3$, and thus $a_{\lambda}b(X_{2})\subset K_{2}$. Finally,
$ba_{\lambda}(X_{1})\subset b\{\operatorname{Re}z>0\}=\{\lvert z-1/4\rvert<1/4\}\subset K_{1}.$
Theorem 1.2 with $\mu=2$ is now an immediate consequence of Corollary 2.10 with $\lambda_{0}=2$.
Let us explain why given $\lambda_{n}\to 2$, we can compute the constants $k_{n}$ mentioned in the theorem effectively. To find a lower bound on the length of the exact subword $w_{n}$ in the theorem, one needs to follow the proofs, which involve determining whether $w(K_{1})$ or $w(K_{2})$ intersect $X_{1}$ and $X_{2}$, for increasingly longer words $w\in\mathcal{B}$ (see Equation (2.4)), until $w\notin W_{K\to X}$. Since the boundaries of these subsets in $\widehat{\mathbb{C}}$ are either lines or circles, this requires computing the image of a finite collection of points under $w$. We also note that our procedure is guaranteed to terminate only when $\lambda_{n}$ corresponds to a non-free group (since the existence of $w\in\mathcal{B}\setminus W_{K\to X}$ arises from the presence of a relator). Therefore, to ensure that $k_{n}$ is computed after finitely many steps, the procedure should be stopped after a prescribed number of iterations, which depends on $\lambda_{n}$ and increases as $n\to\infty$ (so that $k_{n}\to\infty$). ∎
We postpone the proof of the second part of the theorem (with $\mu=i$), and turn to prove Theorem 1.4. Fix some real $21.8\leq\lambda<$ 2 and define $a_{\lambda}$ and $b$ as in Equation (3.1). Then $a_{\lambda},b\in\operatorname{SL}_{2}(\mathbb{R})$, which acts (by Möbius transformations) on ${}\mathbb{R}\cup\{\infty\}$ (the real projective line). We take the intersection of ${}\mathbb{R}\cup\{\infty\}$ with $X_{1},X_{2},K_{1}$ and $K_{2}$ defined in Equations (3.2) and (3.5) and keep the same notation for these subsets, namely
$X_{1}=\{x\in\mathbb{R}:-1<x<1\},$ $X_{2}$ =$=\{x\in\mathbb{R}:\lvert x\rvert>1\}\cup\{\infty\},$
$K_{1}=\{x\in\mathbb{R}:-\frac{1}{2}\leq x\leq\frac{1}{2}\},$ $K_{2}$ =$=\{x\in\mathbb{R}:\lvert x\rvert\geq\frac{5}{4}\}\cup\{\infty\}.$
We wish to understand the set $W_{K\to X}^{\lambda}$, in order to use Corollary 2.7.
Lemma 3.2.
Choose $40<\theta<\pi/$ 4 such that $\lambda=2\cos^{2}\theta$. Then
(3.6) $(a_{\lambda}b^{-1})^{k}K_{2}\subset X_{2}\text{ for every integer }0\leq k\leq\frac{1}{4\theta}.$
Proof.
It will be beneficial to use geometric considerations arising from the action of $\operatorname{SL}_{2}(\mathbb{R})$ on the hyperbolic half-plane $\mathbb{H}^{2}=\{z\in\mathbb{C}:\operatorname{Im}z>0\}$. Denote $c_{\lambda}=a_{\lambda}b^{-1}$. As $\lvert\operatorname{tr}c_{\lambda}\rvert<2$, $c_{\lambda}$ is elliptic and hence acts as a rotation about some point in $\mathbb{H}^{2}$. We recall that a hyperbolic rotation of angle $2\alpha$ about $z\in\mathbb{H}^{2}$ is represented by
(3.7) ±$\pm\begin{pmatrix}\overline{z}&z\\ 1&1\\ \end{pmatrix}\begin{pmatrix}e^{i\alpha}&\\ &e^{-i\alpha}\\ \end{pmatrix}\begin{pmatrix}\overline{z}&z\\ 1&1\\ \end{pmatrix}^{-1},$
and a simple calculation reveals that $c_{\lambda}$ is a rotation of angle $-4\theta$ about $z_{\theta}\coloneqq\frac{1+e^{2i\theta}}{2}$. We claim that (3.6) holds if and only if
(3.8) $c_{\lambda}^{k}(5/4)\in X_{2}\text{ for every integer }0\leq k\leq\frac{1}{4\theta}.$
Since $5/4\in K_{2}$, (3.6) immediately implies (3.8). The other direction should be intuitively clear as well: since $c_{\lambda}$ is a clockwise rotation, the first point of $K_{2}$ to hit $X_{2}^{c}=\{-1\leq x\leq 1\}$ when iterating $c_{\lambda}$ should be the leftmost positive point in $K_{2}$, which is $5/4$; we just need to verify that $5/4$ does not “jump over” $X_{2}^{c}$ when rotated. For $x,y\in\overline{\mathbb{H}^{2}}$, we denote by $[x,y]$ the hyperbolic geodesic segment between $x$ and $y$. Let $\beta$ be the angle between $[z_{\theta},1]$ and $[z_{\theta},-1]$ in $\mathbb{H}^{2}$, measured clockwise; see Figure 1.
Figure 1. The angle between the hyperbolic geodesics $[z_{\theta},1]$ and $[z_{\theta},-1]$, denoted $\beta$, is greater than $\pi$.
Observe that $z_{\theta}$ lies on the geodesic $[0,1]$. Consequently, $\pi<\beta<2\pi$, and in particular, $4\theta<\beta$ which shows the equivalence of (3.6) and (3.8).
Let $\alpha$ be the angle between the geodesics $[z_{\theta},5/4]$ and $[z_{\theta},1]$, measured clockwise (Figure 2).
Figure 2. The angle between $[z_{\theta},\frac{5}{4}]$ and $[z_{\theta},1]$, denoted $\alpha$, is greater than $1$.
Then $\alpha>1$; one way to see it is as follows: for $x\in\mathbb{R}$, let $f_{\theta}(x)$ denote the (hyperbolic) rotation of $x$ by an angle of $1$ about the point $z_{\theta}$. Using Equation (3.7) for $f_{\theta}$, we see that
$f_{\theta}(1)=\frac{\cos(0.5)\cos(\theta)}{\cos(0.5+\theta)}.$
The maximum of $f_{\theta}(1)$ over all $40<\theta^{\prime}<\pi/$ 4 such that $\lambda^{\prime}=2\cos^{2}\theta^{\prime}\geq 1.8$ is attained on $\theta_{0}$ which corresponds to $\lambda_{0}=1.8$, and $f_{\theta_{0}}(1)\approx 1.22$. Thus, $f_{\theta}(1)<5/4$, implying that $\alpha>1$. Therefore, for every integer $0\leq k\leq\frac{1}{4\theta}$,
$4\theta\cdot k\leq 1<\alpha,$
which means that $c_{\lambda}^{k}(5/4)\in X_{2}$ as required. ∎
Proof of Theorem 1.4.
Let $21.8\leq\lambda<$ 2, and choose $40<\theta<\pi/$ 4 such that $\lambda=2\cos^{2}(\theta)$. Let $0\leq k\leq\frac{1}{4\theta}$, $k\in\mathbb{N}$, and denote $u_{\lambda}=(a_{\lambda}b^{-1})^{k}$. By Lemma 3.2, $u_{\lambda}K_{2}\subset X_{2}$. Observe that $a_{\lambda}K_{1}\subset K_{2}$, and recall from the proof of Theorem 1.4 that $b^{-1}X_{2}\subset X_{1}$. Hence, we obtain
$b^{-1}u_{\lambda}K_{2}$ $\subset b^{-1}X_{2}\subset X_{1},$
$u_{\lambda}a_{\lambda}K_{1}$ $\subset u_{\lambda}K_{2}\subset X_{2},$
$(b^{-1}a_{\lambda})^{k+1}K_{1}$ $=b^{-1}u_{\lambda}a_{\lambda}K_{1}\subset X_{1}.$
In other words,
$\{w\in\mathcal{B}:\operatorname{syl}(w)\leq\frac{1}{2\theta}\}\subset W_{K\to X}^{\lambda},$
where $\mathcal{B}$ is the subset from Equation (2.4). By Lemma 3.1, every $w\in\mathcal{B}^{-1}$ with $\operatorname{syl}(w)\leq\frac{1}{2\theta}$ belongs to $W_{K\to X}^{\lambda}$ as well. It follows from Corollary 2.7 that every relator of $H_{\lambda}=G_{\lambda,2}$ has an exact subword from $\mathcal{B}\cup\mathcal{B}^{-1}$ with more than $\frac{1}{2\theta}$ syllables. By possibly omitting the first and last letters, we obtain an exact subword of the desired form. Now an elementary argument shows that
(3.9) $0<\lim_{\lambda\to 2}\frac{\theta}{\sqrt{2-\lambda}}=\lim_{\lambda\to 2}\frac{\arccos{\sqrt{\lambda/2}}}{\sqrt{2-\lambda}}<\infty,$
completing the proof. ∎
Remark 3.3.
The lower bound in Theorem 1.4 cannot be improved without further restrictions on $\lambda$: for every integer $n>2$, choose $\theta_{n}=\pi/2n$. As before, let $\lambda_{n}=2\cos^{2}(\theta_{n})$ and $c_{\lambda_{n}}=a_{\lambda_{n}}b^{-1}$. Then $c_{\lambda_{n}}$ is a (hyperbolic) rotation of angle $-4\theta_{n}=-2\pi/n$, and hence $c_{\lambda_{n}}^{2n}$ is a relator with $4n$ syllables, and as in Equation (3.9), $4n\approx\frac{1}{\theta_{n}}\approx\frac{1}{\sqrt{2-\lambda_{n}}}$.
Proof of Theorem 1.2 for $\mu=i$.
Let $\Lambda=\{\lambda\in\mathbb{C}:\operatorname{Re}\lambda>1.5\}$ as before, and define $a,b:\Lambda\to\operatorname{SL}_{2}(\mathbb{C})$ by
(3.10) $a_{\lambda}=\begin{pmatrix}1&\lambda\\ 0&1\end{pmatrix},\qquad b=b_{\lambda}=\begin{pmatrix}1&0\\ i&1\end{pmatrix}.$
For $z\in\mathbb{C}$ and $r>0$ denote $\mathfrak{B}_{r}(z)=\{w\in\mathbb{C}:\lvert z-w\rvert<r\}$, and define
$X_{1}=\mathfrak{B}_{1}(i)\cup\mathfrak{B}_{1}(0)\cup\mathfrak{B}_{1}(-i),\qquad X_{2}=\{z\in\mathbb{C}:\lvert\operatorname{Re}z\rvert>1\}\cup\{\lvert z\rvert>2\}\cup\{\infty\}.$
It is easy to see that $a_{2}^{n}X_{1}\subset X_{2}$ for every $n\neq 0$, and that we can choose some compact $K_{2}\subset X_{2}$ such that $a_{\lambda}^{n}X_{1}\subset K_{2}$ for every $\lvert n\rvert\geq 2$ and $\lambda\in\Lambda$. Following [18, Theorem 2], one simple geometric way to see the inclusions for $b$ is to set $J=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$, $b^{\prime}z=(Jb^{-1}J^{-1})z=z+i$ and
$X_{1}^{\prime}=JX_{1}=\{\lvert\operatorname{Im}\rvert>1/2\}\cup\{\lvert z\rvert>1\}\cup\{\infty\},\qquad X_{2}^{\prime}=JX_{2}=\mathfrak{B}_{\frac{1}{2}}(\frac{1}{2})\cup\mathfrak{B}_{\frac{1}{2}}(0)\cup\mathfrak{B}_{\frac{1}{2}}(-\frac{1}{2}).$
Then once again it is clear that $(b^{\prime})^{n}X_{2}^{\prime}\subset X_{1}^{\prime}$ for every $n\neq 0$, and that we can choose some compact $K_{1}^{\prime}\subset X_{1}^{\prime}$ such that $(b^{\prime})^{n}X_{2}^{\prime}\subset K_{1}^{\prime}$ for every $\lvert n\rvert\geq 2$. The analogous inclusions for $b^{n}$ now follow with $K_{1}=J^{-1}K_{1}^{\prime}$, and we conclude by applying Corollary 2.10. The constants $(k_{n})$ mentioned in the theorem can be computed effectively for reasons similar to those discussed in the proof of the first part of the theorem. ∎
Proof of Corollary 1.3.
Since $G_{\lambda_{n},\mu_{n}}=G_{\lambda_{n},-\mu_{n}}$, by replacing $\mu_{n}$ with $-\mu_{n}$ if necessary we may assume that $\ell\in\{4,2i\}$. For every (sufficiently large) $n$, choose $z_{n}\in\mathbb{C}$ such that $\mu_{n}z_{n}^{-2}=\ell/2$. Then
$J_{n}B_{\mu_{n}}J_{n}^{-1}=B_{\ell/2},\quad J_{n}A_{\lambda_{n}}J_{n}^{-1}=A_{\lambda_{n}^{\prime}},\qquad\text{where }J_{n}=\begin{pmatrix}z_{n}&0\\ 0&z_{n}^{-1}\end{pmatrix}\text{ and }\lambda_{n}^{\prime}=\lambda_{n}z_{n}^{2}.$
Then $\lambda_{n}^{\prime}\to 2$ and the assertion follows from Theorem 1.2 and the fact that $J_{n}$ conjugates the relators of $G_{\lambda_{n},\mu_{n}}$ to those of $G_{\lambda_{n}^{\prime},\ell/2}$. ∎
Acknowledgements: The author is grateful to Arie Levit for many helpful ideas, suggestions and comments that contributed to the development of this paper, and to Nir Lazarovich for his valuable insight.
The author was supported by ISF grant 1788/22.
[…]